OpenStudy (anonymous):

f(g(h(-2))) f(x)=x^2 g(x)=2x-1 h(x)=3x+7

OpenStudy (anonymous):

gh(x)=6x+14-1 fgh(x)=(6x+13)^2 fgh(-2)=-12+13)^2=1^2=1

OpenStudy (anonymous):

h(-2)=-6+7=1 g(1)=2-1=1 f(g(h(-2)))=f(1)=1

OpenStudy (lalaly):

h(-2)=- 6+7=1 g(1)=1 f(1)=1 so f(g(h(-2))=1

OpenStudy (anonymous):

so its 1?

OpenStudy (lalaly):

yehh

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanx